Pattison family considers the opportunity to finance the purchase of their first $500 thousand worth single-family house in Santa Rosa, CA. To qualify for a 30-year fixed-rate mortgage, the down payment should be 10% of the purchase price. To accumulate money for the down payment, Mr. Pattison plans to invest $35,000 in the stock portfolio that earns 7.65% per annum. Assume that Mr. Pattison reinvests annual gains at the same interest rate as the initial investment. How much money will Mr. Pattison accumulate with this investment within five years? Is the initial investment amount sufficient to accumulate in five years the required down payment under the above mortgage? Round your answers to the nearest dollar. Show your work


1
Expert's answer
2021-05-03T10:52:17-0400

the down payment:

"500 000\times0.1=50 000"

dividends on shares:

"35 000\times0.0765=2677.5=2678"

if he puts in an annual dividend:

"FV=2677.5(1+0.0765)^5=3870.79=3871"

then he does not have enough money for the initial payment

"FV=35000(1+0.0765)^5=50598.57=50599"

if he invests 35,000,then he will get a down payment on the initial payment